# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

### v2

@@ 3. DoubleAgents System
In this section, we provide a walkthrough example to illustrate how DoubleAgents operates in practice. We also describe the implementation details of DoubleAgents. Consider the example of Lee, a PhD student organizing the HCI seminar series at his university. Lee plans to invite four speakers–Prof. A (a famous professor), Prof. B (a rising star professor), Student X (a senior PhD student), and Student Y (a PhD student who is a friend of Lee) – to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5.
Lee begins at the landing page, where he specifies his goals to the system. He types: “I am trying to organize a seminar with four speakers and four slots, help me reach out”. He then provides additional details, including the seminar location, and the dates of the four time slots. Next, Lee enters persona information for each speaker he wants to invite, which the system uses to generate simulated speaker profiles. The landing page is shown in Figure 4. To further customize communication, Lee uploads five of his past emails, allowing the system to adapt its generated messages to his style; for instance, casual for friends and formal for faculty. Finally, Lee initiates the simulation by clicking the “Start Planning” button, prompting the system to set up the scenario with the specified speakers and time slots. At this point, he is also shown a list of preset policies (see Figure 3(A)), which serve as guidelines for the coordination agent to begin planning. Lee reviews the policies, all of which appear reasonable. He may modify existing policies or introduce new ones to better guide the agent’s planning process.
3.1. Context Management Module
At the beginning of the scheduling process, all information provided by Lee on the landing page is compiled into a structured problem context that the system consults at each step to solve the assignment task. DoubleAgents represents this problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history, etc.), and each value contains the corresponding content. As shown in Figure 2, an example snapshot of the problem context includes keys such as ”time info”, which reflects the seminar slots specified by Lee. The context is dynamically updated throughout the interaction, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
3.2. Interactive Chat Interface
Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 3(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents. All plans, actions, and emails generated by the system are accessible through this interface. Each system update appears as an interactive pop-up window (Figure 3(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way. The interface also supports natural language queries, giving Lee full control over what to view. For example, he can type “What did Prof. B reply in the last email?” to retrieve the relevant message history directly within the chat. This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix C.4.
(a) The pop-up window for a generated plan.
(b) The pop-up window for generated actions following the plan.
(c) The pop-up window for a generated email.
Figure 5. Examples for plan and action generation.
3.3. Coordination Agent
3.3.1. Progress Understanding
When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far. To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee. Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups. The prompt used in generating the summary is illustrated in the Appendix C.1.1.
3.3.2. Policy Selection
To design a set of general policies applicable to a broad range of planning problems, the lead authors brainstorm and define rules tailored to address common scheduling challenges. These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants. Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans. Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet. As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state. Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 3(A)), providing Lee with the rationale to generate the plan. The policy acts as the planning guideline, informing the agent’s next steps. Details of the prompts used for the policy selection module are provided in the Appendix C.1.2.
3.3.3. Plan Generation
Guided by this policy, the coordination agent generates the following plan (shown in Figure 5(a)) for Lee: “Send email to Prof. A, Prof. B, Student X, and Student Y to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”. Lee considers the plan and the rationale logical, since scheduling cannot proceed without first collecting availability information, and clicks ’Approve & Continue’ to move forward. The plan and its rationale are automatically recorded in the interactive chat interface (shown in Figure 2(B)), allowing Lee to review past activities and understand the rationale behind the plan. If Lee is dissatisfied with the proposed plan, he can either request the system to regenerate an alternative or directly edit the current plan himself, as shown in Figure 5(a).
3.3.4. Action Generation
To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 5(b)). Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review. He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed. Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input. In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents. The detailed prompts used for plan and action generation are provided in the Appendix C.1.3.
3.3.5. Action Execution
Following approval, the coordination agent proceeds to instantiate each action by generating email drafts for the designated speakers, including both subject lines and message bodies (Figure 5(c)). This is done by calling the email writing tool, as specified in the tool box shown in Figure 2. Specifically, it calls the writeEmail() API, which drafts emails given the current problem state, seminar details, user-provided templates, and historical email edits. These elements allow the system to learn and adapt to the Lee’s preferred communication style, capturing tone, level of formality, and domain-specific phrasing, when generating new emails. Note that email drafting is performed using GPT‑4o, which is invoked by the coordination agent during the execution of the writeEmail() tool.
Lee reviews each draft and identifies one instance where the tone is not appropriate—the email to Prof. A, a distinguished professor, feels overly informal. To address this, Lee types: “Please make the email sound more formal, as Prof. A is a distinguished professor I’d like to invite.” The system then regenerates a more suitable version. Lee further refines the message by editing the first paragraph to better reflect his personal voice: “My name is Lee, a PhD student in Computer Science. Your work in [XXX] has greatly inspired us, and I would be honored to invite you to speak at our HCI seminar series.” After reviewing and finalizing all drafts, Lee is satisfied and clicks ‘Approve & Send’, prompting the system to dispatch the emails to all invited speakers. This is accomplished by invoking the sendEmail() API by the coordination agent, targeting the appropriate speaker email addresses specified in the problem context. In addition, all email-related activities are logged in the dashboard (Figure 2(E)), enabling Lee to easily monitor the status and history of interactions with each speaker. Besides writing and sending emails, the coordination agent may invoke actions involving delayed responses, such as waiting for a speaker’s reply. This is achieved by calling the waitResponse() tool. In such cases, no message is generated or sent; instead, the agent pauses execution and either proceeds to the next pending action or advances to the next simulated day. Overall, we use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
3.3.6. Implicit User Preference Learning
DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful. The system treats such edits as preference signals. To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context. This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time. To streamline the summarization process, we leverage GPT-4o. Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix C.1.5. The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history. This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
3.4. Simulated Respondent Agent
To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics. For instance, the agent representing Prof. A opts not to respond on the same day (shown in Figure 2), as his persona specifies: “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond.” In contrast, Prof. B’s persona indicates she is prompt in communication, leading her agent to reply immediately. Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option. The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix C.3. To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end. Figure 6(a) shows an example response received by Lee as shown in the interface. These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps. Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions. GPT-4o is used when invoking the sendEmail() tool to generate email responses.
(a) Email sent by the simulated Prof. A.
3.5. Edge Case Detector
Within a day, Lee receives an email from Prof. B stating: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person? That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee. Through the Interactive Chat Interface, Lee sees the following system message: “Prof. B requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof. B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof. B accordingly. Lee reviews the suggested action, finds it appropriate, and clicks Approve. An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof. B. Lee’s clarification is recorded as a stop hook and used to refine the policy set, explicitly specifying that Zoom is not preferred when speakers request remote participation. In this way, user interventions during edge cases are converted into reusable system policies. In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans. When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed. The edge case detector is implemented with GPT-4o (see prompts in the Appendix C.2).
3.6. Dashboard Visualization
To support user understanding and foster alignment, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 3(D), (E), and (F), which we describe in detail below:
  • Assignment Grid and Progress Tracker (Figure 3D): This panel visualizes tentative and confirmed speaker-to-slot assignments. It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled. This layer helps Lee assess how close the current plan is to completion at a glance.
  • Calendar and Timelines (Figure 3F): Anchored in simulated days, this temporal layer helps Lee reason about timing, deadlines, and follow-up actions. The calendar highlights both the current simulated day and the seminar dates, allowing Lee to prioritize actions based on urgency and scheduling constraints.
  • Communication Visualization (Figure 3E): This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status. It enables Lee to track ongoing communications, detect gaps, and assess follow-up needs.
These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time. Throughout the process, Lee interacts with the coordination agent and the respondent agents to collaboratively establish speaker-to-slot assignments for the seminar. Together, these visualization features provide Lee with a clear overview of assignment progress and a sense of what needs to be done on the current and subsequent days.
At the end of each simulated day, the system generates a daily summary that highlights key events and decisions. Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix C.5. As shown in Figure 3(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs. A, B, and Students X, Y. Prof. B and Student Y expressed interest, with scheduling pending. Both Prof. B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed. This iterative, day-by-day process continues until all seminar slots are successfully assigned.
@@ 4. Technical Evaluation
DoubleAgents enables users to engage in iterative task planning and action execution through two core functionalities: (1) policy selection, which guides the generation of plans aligned with predefined coordination strategies; and (2) robust edge case detection, which identifies problematic or ambiguous situations and escalates them to the user for clarification and resolution. In this section, we describe the technical evaluation of the two components.
@@ 4.1.1. Experiment Settings
Policy selection refers to the process of identifying relevant coordination policies that guide the generation of high-quality, context-appropriate plans. Specifically, given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module is expected to retrieve all applicable policies that can serve as guidelines for the next planning step. For example, if no speakers have been contacted yet, a suitable policy would be: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” In our implementation, we adopt a seed set of general-purpose coordination policies curated by domain experts. To inform policy selection, we use the output of the getProgressSummary() API, which provides a high-level snapshot of the current task state. An example summary reads: “One slot has been secured: speaker 4 confirmed for Nov 7 (2025-11-07) and received a confirmation email on 2025-08-07. All other speakers are still pending. Last outreach dates: speaker 1 and speaker 2 were last emailed on 2025-08-08; speaker 3 was last emailed on 2025-08-07.
We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs. unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries. Prompt templates for all conditions are provided in Appendix D. To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation. All evaluations were conducted using the o4-mini model with a default temperature of 1. We report the average F1 score across three independent runs for each experimental setting.

### v3

@@ 3.1. User Walkthrough
To illustrate how DoubleAgents operates in practice, we walk through a concrete example. Consider Lee, a PhD student organizing the HCI seminar series at his university. Lee plans to invite four speakers, Prof. Adrian Smith (a famous professor), Prof. Lucy Tran (a rising star professor), Michael Chen (a senior PhD student), and Sarah Malik (a PhD student who is a friend of Lee) to deliver talks during four available time slots: Sep 5, Oct 4, Nov 7, and Dec 5. We demonstrate three example cases of how Lee uses DoubleAgents to solve the problem: (a) starting email bunch, (b) handling edge cases, and (c) following up with non-responsive speakers.
3.1.1. Setting up
Lee begins at the landing page (Figure 6), where he specifies his goal: “I am trying to organize a seminar with four speakers and four slots—help me reach out.” He then provides additional details, including the seminar location and the dates of the four time slots. Next, Lee enters information for each speaker, such as their names and email addresses. When operating in a simulated setting, Lee can optionally include brief persona descriptions to approximate the behavior of the speaker. For example, Prof. Adrian Smith is described as “a famous professor who is very busy and often does not reply to emails,” while Sarah Malik is “a friend who is generally responsive.”. In real-world deployment, these personas are not required, as interactions occur with actual participants. Finally, Lee uploads several past emails he has sent to seminar speakers. This allows the system to adapt to his communication style, generating messages that are more casual for friends and more formal for faculty.
With this information, the system could send emails to speakers on Lee’s behalf. But Lee is nervous about this. He wants to make sure the system sends well-written, personalized emails; follows up appropriately if speakers do not respond; nudges people without being pushy; fills all four slots without double-booking; and, critically, does not offend a famous faculty member with an overly casual message. Lee does not yet know how the system will behave, so rather than letting it act in the real world, he starts a simulation that will not send any actual emails. Instead, the simulation lets him see exactly how the system reasons, what plans it proposes, and how it responds to different speaker behaviors, all guided by a set of policies he can inspect and refine. Lee is also shown a list of preset policies in the Policy Panel (Figure 2(Left)). These are natural-language rules that will guide the AI’s reasoning: for example, when to follow up, how to handle conflicts, and when to give up on a non-responsive speaker. Lee reviews them, finds them reasonable, and is now ready to start the simulation.
3.1.2. DoubleAgents starts and policy selection.
When the simulation starts, Lee sees the Control Panel (Figure 2(Middle)) alongside a dashboard showing the assignment progress, calendar, and communication history (Figure 2(Right)). The calendar shows that in the simulation, today is August 5, and the first seminar date to fill is September 5, so he should start reaching out now. Obviously, nobody has been contacted yet. Lee does not have a clear strategy for who to email first or what to say, so he lets the AI suggest a plan. The coordination agent assesses the current state: no speakers emailed, no availability collected, four slots to fill, and looks for a relevant policy in the existing policy bank, which contains “If someone hasn’t responded to the email after three reach-outs, then don’t contact them again.”, “Don’t be shy to ask for more availability if they don’t give you many. They will probably reply more if you beg.” and a few others. Then the agent finds one that’s most relevant: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is clearly highlighted to Lee in the Policy Panel (Figure 2(Left)), so Lee can see exactly why the AI is about to do what it does. Based on this policy, the AI subsequently proposes a plan (Figure 7(a)): “Send email to Prof. Adrian Smith, Prof. Lucy Tran, Michael Chen, and Sarah Malik to request availability for slots on Sep 5, Oct 4, Nov 7, and Dec 5.”
This makes sense to Lee: scheduling cannot proceed without first collecting availability. So he clicks ‘Approve & Continue.’ The system then generates four concrete actions, one email per speaker (Figure 7(b)), which Lee reviews and approves. Now the AI drafts the actual emails (Figure 7(c)). Lee reads the first draft carefully. He notices the date format is inconsistent and wants to add a note mentioning that his advisor recommended the speaker. He makes these edits and clicks ‘Regenerate.’ The revised draft incorporates all his changes. For the next email to Prof. Lucy Tran, the system has already learned Lee’s preferred date format and applies it automatically, requiring no further corrections. Lee approves and sends all four emails. Even though these are simulated, Lee is reassured: the AI is learning his style, and he can see exactly what would be sent in the real world.
3.1.3. Handling responses from speakers.
In reality, Lee would have to wait days or weeks for responses. In the simulation, time is compressed: a full day in the real world takes only one to two minutes. So Lee quickly sees what happens next. By the end of Day 1, three of the four speakers have responded. Michael Chen and Sarah Malik reply with their availability. Prof. Lucy Tran also replies promptly, but with an unexpected request: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person?” Prof. Adrian Smith, true to his persona, does not respond at all. The assignment progress tracker immediately updates to reflect who has replied and what slots they are available for. The communication log records every exchange. Lee can see at a glance where things stand and where the gaps are.
3.1.4. Edge cases handling.
Prof. Lucy Tran’s Zoom request creates a problem: none of the existing policies address remote presentations. The AI does not try to handle this on its own. Instead, the edge case detector flags it and asks Lee for guidance: “Prof. Lucy Tran requests a Zoom option. Should I confirm Zoom availability or insist on in-person participation?” This is exactly the kind of situation Lee was worried about: the AI encountering something unexpected and making a bad call. But instead of acting autonomously, the system pauses and asks. Lee decides that Zoom is acceptable and types: “Zoom is allowed.” The AI generates a response to Prof. Lucy Tran, which Lee reviews and sends. Importantly, Lee’s decision does not resolve this case alone. The system records it as a stop hook and updates the policy set to include a new rule about allowing Zoom when speakers request it. If a similar situation arises later, for example, Michael Chen also asks about Zoom, the system will already know how to handle it without bothering Lee again. At the end of Day 1, the system generates a daily summary of everything that happened: which emails were sent, who replied, and what decisions were made. Lee reviews it, finds it accurate, and clicks ‘Continue to Next Day.’
3.1.5. Clarifying, waiting, and following up.
The calendar advances. Prof. Adrian Smith has not yet responded, while Prof. Lucy Tran’s Zoom details remain unresolved. The AI generates a plan: clarify Zoom logistics with Prof. Tran and continue waiting for Prof. Smith (Figure 2(Middle)). Lee checks the rationale, confirms it references the appropriate policies, and approves. The system drafts a follow-up to Prof. Tran, who replies confirming November 7th via Zoom. The dashboard updates accordingly, leaving October 4th as the only open slot.
On Day 3, Prof. Smith has still not responded. The coordination agent, guided by the follow-up policy specifying action after two workdays of silence, drafts a gentle nudge. Lee notices it matches his usual follow-up style: short and to the point, with no unnecessary pressure. This is something the system picked up from uploaded emails and his earlier edits. He approves without changes. Later on, Prof. Smith replies with availability on November 7th, creating a new conflict: three speakers now favor the same date while October 4th remains unfilled. Lee can see the conflict clearly in the assignment progress tracker. From this point onward, the system continues to generate plans grounded in policies, learn from Lee’s past choices, and execute actions adaptively, resolving the scheduling conflict and filling the remaining slot until the seminar assignment is complete.
3.1.6. Usage conclusion.
By the time the DoubleAgents usage concludes, Lee has not only produced a tentative speaker-to-slot assignment but has also built up a set of refined policies, customized email templates, and executable stop hooks that align with his actual preferences. He now knows how the AI handles follow-ups, edge cases, and conflicts, because he has seen it happen in a safe and simulated environment. When moving to real-world deployment, such as sending actual emails to speakers, these artifacts directly guide the agent’s actions, reducing uncertainty and enabling more reliable coordination. Over time, this process accumulates into a reusable set of alignment artifacts that can be applied and further developed in future coordination tasks.
3.2. System Implementations
We implemented DoubleAgents as an interactive simulation system built on the ReAct framework (Yao et al., 2023). The system architecture comprises four key components: (1) a Simulation Engine that drives an iterative planning-execution loop through a coordination agent and LLM-based simulated respondent agents, (2) a Policy Module that encodes user values as natural language rules, supports policy selection and refinement through stop hooks, and captures implicit user preferences from edit behavior, (3) an Edge Case Detector that identifies situations beyond existing policy coverage and escalates them for user clarification, and (4) a Dashboard Visualization built on Flask that provides real-time visual feedback on assignment progress, communication history, and scheduling timelines. Users interact with the system via a control panel that supports plan review, action editing, and natural-language queries. Full implementation details, including prompts and system architecture, are provided in Appendix B.
@@ 4. Technical Evaluation
DoubleAgents enables iterative task planning and action execution through two core mechanisms: (1) policy selection, which steers plan generation toward user-aligned coordination strategies; and (2) edge case detection, which identifies ambiguous or potentially problematic situations and escalates them for user intervention. Together, these components support a structured interaction loop in which users can guide, refine, and correct agent behavior over time. We present a technical evaluation of these two components.
@@ 4.1.1. Experiment Settings
Policy selection refers to identifying coordination policies that guide the generation of high-quality, user-aligned plans. Given a particular problem state (e.g., unfilled time slots and unresponsive speakers), the policy selection module retrieves relevant policies that can inform the next planning step. For example, if no speakers have been contacted yet, a suitable policy would be: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” In our implementation, we adopt a seed set of general-purpose coordination policies curated by domain experts.
To study how different representations of the problem state affect policy selection, we compare four input conditions that vary in the level of abstraction and context provided to the model:
  • Speaker-only context, which includes information about each speaker’s persona. For example: “Speaker 1 (Prof. Adrian Smith): busy, often replies only after multiple follow-ups; Speaker 2 (Dr. Lucy Tran): usually responds within 1–2 days; Speaker 3 (Sarah Malik): highly responsive and flexible; Speaker 4 (Michael Chen): prefers early scheduling but has not replied recently.”.
  • TimeSlot-only context, which includes the current status of all time slots and their constraints. For example: “Four slots: Sep 5 (unfilled), Oct 3 (filled), Nov 7 (filled), Dec 5 (unfilled). The Nov 7 slot has been confirmed with a speaker, while Oct 3 is tentatively held. Remaining open slots require confirmation.
  • Combined TimeSlot and Speaker context, which includes both sources of information without additional abstraction. For example: “Speaker 1 (busy, slow to reply) and Speaker 2 (moderately responsive) were last contacted two days ago; Speaker 3 (responsive) has not yet been contacted; Speaker 4 prefers early slots but has not replied. Slot status: Oct 3 and Nov 7 are filled (Nov 7 confirmed), while Sep 5 and Dec 5 remain open.
  • Summary-based context, which includes the information of the speakers, time slots, and the summary of the current task state to provide a structured, high-level snapshot of the system. An example summary reads: “One slot has been secured: speaker 4 confirmed for Nov 7 (2025-11-07) and received a confirmation email on 2025-08-07. All other speakers are still pending. Last outreach dates: speaker 1 and speaker 2 were last emailed on 2025-08-08; speaker 3 was last emailed on 2025-08-07.
Prompt templates for all conditions are provided in Appendix E. We construct a test suite of 30 examples sourced from real user study sessions, covering a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as ground truth. All evaluations are conducted using the o4-mini model with a default temperature of 1, and we report the average F1 score across three independent runs for each condition.